OpenStudy (anonymous):

find the vertical, horizontal, or oblique asymptote for the equation. r(x)=x^2+x-6/x-3

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

The vertical asymptotes occur for values of x that make the rational function undefined (i.e. the values of x that make the denominator = 0). In this case, values of x for which the following holds: \(x-3=0\) That is, assuming your function is \[f(x)=\frac{x^2+x-6}{x-3}\] Horizontal asymptotes occur at the values of y that are equal to the limits at infinity; in other words, a horizontal asymptote occurs at the lines \(y=L_1\text{ and }y=L_2\) if \[\lim_{x\to\infty}f(x)=L_1,\text{ and }\lim_{x\to-\infty}f(x)=L_2\] If these limits are not finite, the function has oblique/slant asymptotes that take on the form of the remainder of the quotient. That is, using long division, the oblique asymptote will be \[y=\frac{R(x)}{x-3},\text{ where }R(x)\text{ is the remainder}.\]